OpenStudy (anonymous):

the density of an object is calculated using the formula D=m/v, where m is the objects mass and v is its volume. Gold has a density of 19.3g/cm^3. What is the volume of an amount of gold that has a mass of 96.5g?

OpenStudy (anonymous):

D = m/v isolate v v = m/D, solve for v by plugging in your known variables.

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

Your welcome.